Circular No. 6136

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

A0538–66 R. Corbet and A. Smale, Goddard Space Flight Center and Universi-

ties Space Research Association; and P. Charles and K. Southwell, Oxford University, write: “The recurrent x-ray transient A0538–66 (16.65-day period) remains active although at a low level. ASCA observations on Feb. 3 (predicted x-ray maximum) and 4 gave GIS count rates of ∼ 0.1 and 0.05 counts/s, respectively. The estimated luminosity of 2 × 10^{36} erg/s is much less than the largest outbursts previously observed. Photometry from the South African Astronomical Observatory yielded the following (conservative errors of 0.03 mag): Jan. 31, V = 15.05; Feb. 1, 14.67; 2, 14.99; 3, 14.86; 4, 14.98; 5, 15.04; 6, 15.04. Spectra obtained at the Anglo-Australian Telescope on 1994 Nov. 29, shortly after a previous predicted outburst, showed variable Balmer emission but not the He II 468.6-nm emission that accompanies large outbursts. Optical observations around future predicted maxima are encouraged.”

COMET C/1994 T1 (MACHHOLZ) Ephemeris continuation (cf. IAUC 6113), from orbital elements on

MPC 24711:

1995 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Feb. 12 2^{h}58.^{m}16 − 6^{o}23.^{′}4 2.467 2.462 78.^{o}1 23.^{o}1 13.8
17 3 00.13 − 6 58.0 2.588 2.501 73.9 22.3 14.0
22 3 02.51 − 7 27.8 2.707 2.541 69.8 21.4 14.2
27 3 05.24 − 7 53.9 2.822 2.581 65.9 20.5 14.4
Mar. 4 3 08.27 − 8 17.1 2.934 2.621 62.2 19.5 14.6
9 3 11.57 − 8 38.1 3.042 2.663 58.6 18.6 14.8
14 3 15.10 − 8 57.7 3.145 2.704 55.2 17.6 15.0
19 3 18.83 − 9 16.1 3.245 2.746 52.0 16.6 15.1
24 3 22.74 − 9 34.0 3.339 2.788 49.0 15.6 15.3
29 3 26.80 − 9 51.7 3.429 2.831 46.1 14.7 15.5
Apr. 3 3 31.00 −10 09.5 3.515 2.873 43.5 13.9 15.6
8 3 35.31 −10 27.8 3.595 2.916 41.2 13.1 15.8
13 3 39.71 −10 46.9 3.670 2.960 39.1 12.3 15.9
18 3 44.20 −11 07.0 3.740 3.003 37.3 11.7 16.0
23 3 48.75 −11 28.2 3.805 3.047 35.8 11.1 16.2

1995 February 16 Daniel W. E. Green